OpenStudy (abmon98):

Suppose that f satisfies the equation f(x + y) = f(x) + f(y) + x^2y + xy^2 for all real numbers x and y.

OpenStudy (abmon98):

Suppose further that \[\lim_{x \rightarrow 0} f(x)/x=1\] a) Find f(0) b) Find f'(0) c) Find f'(x)

OpenStudy (abmon98):

Here is my working f(x+y)=f(0) therefore x=-y So f(x+(-x))=f(x)+f(-y) +0 +0 f(0)=f(x)+f(-x)=0 Is this wrong?

zepdrix (zepdrix):

It looks correct up to this point: f(0) = f(x) + f(-x) not sure why you're saying that =0 though. And I'm not sure if this is the approach they wanted, hmm still thinking about it :)

OpenStudy (abmon98):

\[f \prime(0)=\lim_{h \rightarrow 0} f(0+h)-f(0)/h=\lim_{h \rightarrow 0} f(h)-0/h=1\]

OpenStudy (abmon98):

Can we say that x=0

zepdrix (zepdrix):

\[\large\rm f(\color{orangered}{x}+y)=f(\color{orangered}{x})+f(y)+\color{orangered}{x}^2y+\color{orangered}{x}y^2\]\[\large\rm f(\color{orangered}{0}+y)=f(\color{orangered}{0})+f(y)+\color{orangered}{0}^2y+\color{orangered}{0}y^2\]\[\large\rm f(y)=f(0)+f(y)\]So that's where you're getting f(0)=0, ok ok clever.

zepdrix (zepdrix):

*Subtracting f(y) from each side* just in case that step wasn't clear.

zepdrix (zepdrix):

I dunno if that's right, bahhh we need someone smart

zepdrix (zepdrix):

@ganeshie8 @SithsAndGiggles

OpenStudy (abmon98):

@zepdrix lol, i think the work you added above is right :) but i am not sure of my approach x=-y

OpenStudy (anonymous):

(a) and (b) look right to me. Setting \(x=-y\) yields \[f(0)=f(x)+f(-x)\]and setting \(x=0\) yields \[f(0)=2f(0)\implies f(0)=0\] For the derivative at \(0\), that follows immediately from the definition of the derivative and the given behavior of \(\dfrac{f(x)}{x}\) near \(x=0\) : \[f'(c)=\lim_{x\to c}\frac{f(x)-f(c)}{x-c}\implies f'(0)=\lim_{x\to0}\frac{f(x)}{x}=1\]

OpenStudy (abmon98):

Can i ask another question is the limit as x approaches 0 of f(x)/x=1 is that the derivative at f'(0)

OpenStudy (anonymous):

Yes

OpenStudy (abmon98):

Thank you @SithsAndGiggles and @zepdrix for your help :)

OpenStudy (abmon98):

f'(x)=lim as h-->0 f(x+h)-f(x)/h =lim as h-->0 f(x)+f(h)+x^2h+xh^2-f(x)/h =1+lim h-->0 hx(x+h)/h=1+x^2

OpenStudy (anonymous):

Yes, that's reasonable.